OpenStudy (anonymous):

Find the square roots of 2500. What is the principal square root of 2500?

OpenStudy (anonymous):

50

jhonyy9 (jhonyy9):

50

OpenStudy (anonymous):

postive or negative 50

OpenStudy (asnaseer):

the square root of a positive number is always plus/minus some other number. e.g.:\[\sqrt{25}=\pm5\]

jhonyy9 (jhonyy9):

every two

OpenStudy (anonymous):

It should be positive since it's asking the principle of square root...

OpenStudy (asnaseer):

so try and break up your number into components you can handle. e.g. 2500 = 25 * 100 and use the fact that:\[\sqrt{ab}=\sqrt{a}*\sqrt{b}\]

OpenStudy (anonymous):

ok What is the negative square root of 2500?

OpenStudy (asnaseer):

@Darkmirror - the first part of the question states "Find the square roots..." - which means find both (I think)

OpenStudy (anonymous):

@asnaseer ~ yup ~joyce ~ the negative one is -50

OpenStudy (anonymous):

THX

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

it depends where u are,here the square root of a positive number is always positive,the negative sqrt is simply -sqrt(of the number)

OpenStudy (asnaseer):

@ghass1978 - where is "here"?

OpenStudy (anonymous):

well french system..

OpenStudy (asnaseer):

and I was always under the impression that Maths was a universal truth.... oh wel....

OpenStudy (anonymous):

:)

OpenStudy (asnaseer):

at least if you had said "in a parallel Universe to where you are" - then I would have understood :-)